OpenStudy (anonymous):

Approximate the solution to the heat equation du/dt-d2u/dx2=0, 0

OpenStudy (anonymous):

Approximate the solution to the heat equation \[\frac{ \delta u }{ \delta t }-\frac{ \delta^{2}u }{ \delta x^{2} }=0\] \[0<t <1, 0 \le t\] with boundary condition \[u(0,t)=u(1,t)=0,0<t\] and initial condition \[u(x,0)=\sin(\pi x)\] Compare results at time t=0.5 at location x=0.2,0.4,0.6,0.8 to the exact solution \[u(x,t)=\exp (-\pi^{2} t)\sin(\pi x)\]